OpenStudy (anonymous):

What is the y-intercept of the function y = 1.2(3.8)x?

OpenStudy (unklerhaukus):

the y intercept is when x=0

OpenStudy (anonymous):

is it 1.2

OpenStudy (unklerhaukus):

\[y(x)= 1.2(3.8)^x\]\[y(0)= 1.2(3.8)^0\]\[y(0)= 1.2\times1=1.2\color{red}\checkmark\]

OpenStudy (anonymous):

so i was correct

OpenStudy (unklerhaukus):

yeah great work

OpenStudy (anonymous):

wooooooooooo